graph the system of inequalities
x + 2y < 6
2x + y < 6
x > 0
y > 0

Answer by Edwin McCravy(20055)


x + 2y < 6
2x + y < 6
     x > 0
     y > 0

Plot the boundary lines whose equations are
like the above inequalities except that the
inequality signs are replaced by equal signs.

x + 2y = 6
2x + y = 6
     x = 0 (the y-axis)
     y = 0 (the x-axis)

  

Shade the part that is under the red and green lines,
and above the x-axis and to the right of the y-axis.

That is, you will end up with just this quadrilateral:


Edwin